OpenStudy (xapproachesinfinity):

@SolomonZelman here is an integral for you lol \(\large \int \frac{2x^2+2x-6}{x-6}dx\) if you still like them hehe

OpenStudy (solomonzelman):

Hey how a bout some magic zeros? \[\Large \int\limits_{ }^{ } \frac{2x^2-2x-6}{x-6}~dx\]\[\Large 2 \int\limits_{ }^{ } \frac{x^2-x-3}{x-6}~dx\]\[\Large 2 \int\limits_{ }^{ } \frac{x^2-x-3-27+27}{x-6}~dx\] \[\Large 2 \int\limits_{ }^{ } \frac{x^2-x-30+27}{x-6}~dx\]\[\Large 2 \int\limits_{ }^{ } \frac{x^2-x-30}{x-6}+\frac{27}{x-6}~dx\]

OpenStudy (solomonzelman):

:PPPPPPP

OpenStudy (solomonzelman):

ha ha and factor the top and on...

OpenStudy (xapproachesinfinity):

hmm interesting

OpenStudy (xapproachesinfinity):

i did it differently i first separated it into 3 integrals then i used u sub u=x-6

OpenStudy (solomonzelman):

The main thing is that it's easy, not that's it's interesting, although it's certainly is:)

OpenStudy (solomonzelman):

my way is easier :PP lol

OpenStudy (xapproachesinfinity):

i guess so^_^

OpenStudy (xapproachesinfinity):

i attempted to do it that way but i didn't see -30 lol

OpenStudy (solomonzelman):

well, i did this on my paper. What gets me x^2 and -x for my middle term with (x-6) ? it was (x+5) .... (x+5)(x-6) the last term is -30.

OpenStudy (xapproachesinfinity):

Oh i see!

OpenStudy (solomonzelman):

it's actually a super helpful thing for lazy ppl.

OpenStudy (xapproachesinfinity):

I'm sometimes lazy lol

OpenStudy (solomonzelman):

and I'm always. well, i mean why make it harder, if it easy in reality?

OpenStudy (solomonzelman):

tnx for throwing a question though, it was good.

OpenStudy (xapproachesinfinity):

lol! my pleasure ^_^

OpenStudy (anonymous):

You really did 3 integrals?

OpenStudy (anonymous):

So you had to integrate \(x^2/(x-6)\)

OpenStudy (solomonzelman):

I didn't have to? why?

OpenStudy (anonymous):

Nothing, I'm just thinking. I guess in the end the \(u=x-6\) is most straight forward approach

OpenStudy (solomonzelman):

u would do u, lol, for me it seemed easy to complete the fraction (using it similarly to "completing the square")

OpenStudy (xapproachesinfinity):

@wio yes that's what i did and one of the integrals is x^2/x-6 with u sub but i think @SolomonZelman is easier

OpenStudy (solomonzelman):

' algebra ' is easier than calc, I would assume so:)

OpenStudy (anonymous):

^ this guy xD